# PBPK-iPINNs: Inverse Physics-Informed Neural Networks for Physiologically Based Pharmacokinetic Brain Models

### v2

@@ Abstract
Physics-Informed Neural Networks (PINNs) leverage machine learning with differential equations to solve direct and inverse problems, ensuring predictions follow physical laws. Physiologically based pharmacokinetic (PBPK) modeling advances beyond classical compartmental approaches by using a mechanistic, physiology focused framework. A PBPK model is based on a system of ODEs, with each equation representing the mass balance of a drug in a compartment, such as an organ or tissue. These ODEs include parameters that reflect physiological, biochemical, and drug-specific characteristics to simulate how the drug moves through the body. In this paper, we introduce PBPK-iPINN, a method to estimate drug-specific or patient-specific parameters and drug concentration profiles in PBPK brain compartment models using inverse PINNs. We demonstrate that, for the inverse problem to converge to the correct solution, the loss function components (data loss, initial conditions loss, and residual loss) must be appropriately weighted, and hyperparameters (including layers, neurons, activation functions, learning rate, optimizer, and collocation points) must be carefully tuned. The performance of the PBPK-iPINN approach is then compared with established traditional numerical and statistical methods. Relevance to Life Sciences: In this study, we consider a permeability-limited, 4-compartment PBPK brain model that mimics human brain functionality in drug delivery. Such models involve many unknown parameters (drug-specific and patient-specific) that are difficult to measure directly in humans due to ethical and practical constraints. We propose inverse PINNs as an alternative approach for estimating these parameters. Accurate parameter estimation yields precise drug concentration–time profiles, which in turn enable the calculation of pharmacokinetic metrics (AUC, Tmax, Cmax) that support drug developers and clinicians in designing and optimizing therapies for brain cancer. Mathematical Content: The aforementioned PBPK models consist of a complex system of linear, nonhomogeneous, parametric ordinary differential equations (ODEs), sometimes with time-dependent parameters. In this study, we present the existence and uniqueness of the forward problem. These ODEs are often stiff, and traditional numerical and statistical methods frequently fail to converge when solving the forward and inverse problems. Therefore, we employ physics-informed neural networks (PINNs) to identify the optimal set of parameters by minimizing the total loss, which includes data losses, ODE residual losses, and initial condition losses.
@@
11footnotetext: Submitted to the editors DATE. 09/12/2025.22footnotetext: Department of Oncology, Wayne State University, Detroit, MI (gi6036@wayne.edu).33footnotetext: Department of Chemistry, Siena Heights University, Adrian, MI (kweerasi@sienaheights.edu).33footnotetext: Department of Engineering, Siena Heights University, Adrian, MI (pranawee@sienaheights.edu).
@@ 1 Introduction
The paper is organized as follows. In section 2 we present the system of ordinary differential equations and the schematic illustration of the 4-compartment brain model that mimic the drug transport in different regions of the human brain and it’s existence and uniqueness of the forward problem. Section 3 is devoted to explain the inverse physics-informed neural network architecture for the four compartment brain model and a step by step guidance to implement a Python code of the PINN algorithm used through DeepXDE library [23]. In Section 4, we present numerical results showing the predictive power of PBPK-iPINN and its validation. Finally, a conclusion is drawn that highlights impact of PBPK-iPINN in Section 5.
@@ 2.1 System of Differential Equations
PSBPSB Passive permeability-surface area of BBB (L/h) 135.0
PSCPSC Passive permeability-surface area of BBB (L/h) 67.5
@@ 4.1 A Step-by-Step Guide to Implementing a Four-Compartment Brain PINN in DeepXDE
    Define exogenous input function: The plasma concentration data are provided as input to the system via the function $g_{1}(t;\theta)$, as defined in Equation (2.6). Since the plasma data are available at discrete time points, linear interpolation is applied to construct a continuous approximation. As a result, $g(\theta,t)$ becomes a continuous function of time. To this end we implemented following function (PlasmaInterp) to our program.
@@ 5 Numerical Results
This section presents our main results and their validation. In Example 1, we identify the optimal configuration including the activation function, number of layers and neurons, weight and bias initializers (Glorot uniform and Glorot normal), optimizer, and learning rate that minimizes the total loss (objective function). In Example 2, we solve the inverse problem, estimating the parameters $V_{\mathrm{bb}}$, $V_{\mathrm{bm}}$, $V_{\mathrm{ccsf}}$, $V_{\mathrm{scsf}}$, $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ and presenting the corresponding predicted concentration profiles. Finally, in Example 3, we validate these results using two additional numerical and statistical techniques.
@@
As shown in Table (3), the Tanh and Sigmoid activation functions generally perform best for the four-compartment brain model. Figure (7) further confirms their accuracy compared to Sin and ReLU by overlaying observed data with the predicted concentration profiles from each activation function. The training times in Table (3) align with the expected trend that computational cost increases with the number of layers and neurons. Based on these results, we select the $\tanh$ activation function with six layers and 50 neurons per layer for further optimization. The entry 4.20e-02 (1017) in the Table (3) indicates a total loss of $4.20\times 10^{-2}$ and a training time of $1017$ seconds for this configuration.
@@
Sin 5.69e-02 (571) 3.16e-02 (573) 4.24e-02 (724) 4.57e-02 (1393)
@@
Example 2. In this example, we successfully estimated up to six selected parameters $V_{\mathrm{bb}}$, $V_{\mathrm{bm}}$, $V_{\mathrm{ccsf}}$, $V_{\mathrm{scsf}}$, $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ though any other parameter set could also be estimated. However, selecting a large number of parameters comes at the cost of increased computational expense. Based on the conclusion of Example 1, the selected activation function is tanh with Glorot normal initializer for initial weight and bias generator. The number of neuron per layer and the number of layers are 50 and 6 respectively. We found that $\lambda_{IC}=(1,1,1,1)$, $\lambda_{ODE}=(2,2,2,2)$, and $\lambda_{Data}=(3,3,3,3)$ yielded the best results in our experiments. Then we trained the network for 5 million iterations, which took 54977.47 seconds to complete. After training the model with adam’s optimizer with the learning rate of 0.0001 we further train the model with Limited-memory Broyden-Fletcher-Goldfarb-Shanno optimizer for further smoothness and the best model is found at the 5000018 iteration with total loss of 1.23e051.23e-05. As evidence in Table (4) it can be seen that the parameters $V_{\mathrm{bb}}$, $V_{\mathrm{bm}}$, $V_{\mathrm{ccsf}}$ and $V_{\mathrm{scsf}}$ were trained with very high accuracy. The estimated parameter values of $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ are physiologically acceptable and accurate enough. However, based on the level of accuracy that we expect the model can be further trained.
@@
Table 4: Comparison of reference parameter values and the model estimated parameter values with absolute errors.
Parameter Reference Value PINN Value Absolute Error
$V_{bb}$ 0.064952435 0.06495242565870285034 $9.34\times 10^{-9}$
$V_{bm}$ 1.104115461 1.10411536693572998047 $9.41\times 10^{-8}$
$V_{ccsf}$ 0.103984624 0.10398460924625396729 $1.48\times 10^{-8}$
$V_{scsf}$ 0.025996156 0.02599614672362804413 $9.28\times 10^{-9}$
$fu_{bb}$ 0.125 0.12873644798994064331 $3.74\times 10^{-3}$
$\lambda_{ccsf}$ 0.026 0.02129559218883514404 $4.70\times 10^{-3}$
@@
Figure 9: Minimization of weighted loss components over epoch.
@@
Example 3. In this example, we compare the prediction accuracy of our PBPK-iPINN approach by comparing it with traditional statistical and numerical methods. First, the Stochastic Approximation Expectation-Maximization (SAEM) algorithm [30], implemented in MONOLIX, was used to estimate model parameters. The forward problem was then solved using explicit Runge-Kutta (RK) schemes. Second, the Differential Evolution (DE) algorithm [31], implemented in R, was used for parameter estimation. The forward problem was subsequently solved using the Livermore Solver for Ordinary Differential Equations with Adaptive step-size control, which automatically switches between a non-stiff method (Adams) and a stiff method (Backward Differentiation Formula, BDF) depending on problem stiffness. Physiologically meaningful parameter bounds were applied across all methods to constrain the parameter space and accelerate convergence. For a fair comparison, identical parameter bounds were defined for the PBPK-iPINN, SAEM, and DE algorithms. Absolute errors (for example, $|\text{PINN}_{\text{err}}|=|\text{Reference Value}-\text{PINN}|$), reported in Table (5), were calculated relative to the reference parameter values and rounded to five decimal places. As summarized in Table (5), PBPK-iPINN demonstrates promising potential as an alternative to traditional numerical and statistical methods for addressing complex inverse problems, particularly those involving highly stiff systems.

### v3

@@ Abstract
Physics-Informed Neural Networks (PINNs) integrate machine learning with differential equations to solve forward and inverse problems while ensuring that predictions adhere to physical laws. Physiologically based pharmacokinetic (PBPK) modeling advances beyond classical compartmental approaches by employing a mechanistic, physiology-focused framework. Such models involve many unknown parameters that are difficult to measure directly in humans due to ethical and practical constraints. PBPK models are constructed as systems of ordinary differential equations (ODEs) and these parametric ODEs are often stiff, and traditional numerical and statistical methods frequently fail to converge. In this study, we consider a permeability-limited, four-compartment PBPK brain model that mimics human brain functionality in drug delivery. We introduce PBPK-iPINN, a method for estimating drug-specific or patient-specific parameters and drug concentration profiles using inverse PINNs. We also conducted parameter identifiability analysis to determines whether the parameters can be uniquely and reliably estimated from the available data. We demonstrate that, for the inverse problem to converge to the correct solution, the components of the loss function (data loss, initial condition loss, and residual loss) must be appropriately weighted, and the hyperparameters including the number of layers and neurons, activation functions, learning rate, optimizer, and collocation points must be carefully tuned. The performance of the PBPK-iPINN approach is then compared with established numerical and statistical methods. PBPK-iPINN achieved parameter errors on the order of $10^{-8}$ for volumes, outperforming traditional methods. Accurate parameter estimation yields precise drug concentration-time profiles, which in turn enable the calculation of pharmacokinetic metrics. These metrics support drug developers and clinicians in designing and optimizing therapies for brain cancer.
@@
22footnotetext: Department of Oncology, Wayne State University, MI (gi6036@wayne.edu).33footnotetext: Department of Chemistry, Siena Heights University, MI (kweerasi@sienaheights.edu).33footnotetext: Department of Engineering, Siena Heights University, MI (pranawee@sienaheights.edu).33footnotetext: Department of Mathematics and Statistics, Northern Kentucky University, KY (hapuhinnan1@nku.edu).
@@ 1 Introduction
Parameter identifiability is a fundamental concept in system modeling that determines whether model parameters can be uniquely estimated from available data. Using the four-compartment brain model, we assess the structural identifiability of the model parameters. When parameters are found to be structurally unidentifiable, we fix selected parameters, leveraging the richness of the available dataset. Subsequently, using the parameters inferred through the PBPK-iPINN framework, we perform a practical identifiability analysis to evaluate the quality and reliability of the parameter estimates
@@
Figure 2: Advancing pharmacokinetic modeling with neural network.
@@
The paper is organized as follows. In section 2 we present the system of ordinary differential equations and the schematic illustration of the 4-compartment brain model that mimic the drug transport in different regions of the human brain and it’s existence and uniqueness of the forward problem. We further conduct structural and practical identifiability analyses to assess whether the model parameters can be uniquely determined. Section 3 is devoted to explain the inverse physics-informed neural network architecture for the four compartment brain model and a step by step guidance to implement a Python code of the PINN algorithm used through DeepXDE library [23]. In Section 4, we present numerical results showing the predictive power of PBPK-iPINN and its validation. Finally, a conclusion is drawn that highlights impact of PBPK-iPINN in Section 5.
@@ 2.1 System of Differential Equations
PSBPSB Passive permeability-surface area product at the BBB (L/h) 135.0
PSCPSC Passive permeability-surface area product at the blood–cranial CSF barrier (L/h) 67.5
@@ 2.2 Existence and Uniqueness
2.3 Parameter Identifiability Analysis
In this section, to complete the system identification workflow, we perform structural and practical identifiability analyses to assess whether the model parameters can be uniquely determined from the available data.
Structural Identifiability Analysis
We consider the following general form of the mathematical model for parameter estimation:
$\quad\dot{x}(t,p)=f(x(t),u(t),p),\quad y(t,p)=g(x(t),p),\quad x_{0}=x(t_{0},p)$ (2.11)
where $f$ and $g$ are vector functions of their arguments, $p\in\mathbb{R}^{p}$ is a $p$-dimensional vector of parameters, $x(t)\in\mathbb{R}^{n}$ is the $n$-dimensional state variable vector, $u(t)\in\mathbb{R}^{r}$ is the $r$-dimensional input vector, and $y(t)\in\mathbb{R}^{m}$ are the $m$-dimensional measured outputs. A parameter $p_{i}$ is said to be identifiable if the following equation holds true:
$y(t,p)=y(t,p^{*})\quad\Rightarrow\quad p_{i}=p_{i}^{*}$ (2.12)
where $p^{*}$ is an alternative parameter vector.
  • If Eq.(2.12) holds for any $p_{i}^{*}$, then $p_{i}$ is said to be globally identifiable.
  • If Eq.(2.12) holds for a neighborhood of $p_{i}^{*}$, then $p_{i}$ is said to be locally identifiable.
  • If Eq.(2.12) does not hold true for any $p_{i}$ locally or globally, then $p_{i}$ is said to be structurally unidentifiable.
There are multiple methods that can be used to perform structural identifiability analysis [36, 37]. In this section, we only test for the local identifiability of the system and, for convenience, we will refer to a system as being identifiable when it is structually locally identifiable. We use the Julia library StructuralIdentifiability [38] to test for structural identifiability of the model. The existing algorithms implemented in the library require both f and g to be rational functions, which are fractions of polynomials. We specify the parametric ODE model using the @ODEmodel macro. x’(t) is the derivative of state variable x(t), which is assumed to be unknown if not specified otherwise. y(t) defines the output variable which is assumed to be given. Figure (4) shows the Julia code to perform local structural identifiability. The last line of the code tests the local identifiability of the model.
Figure 4: Julia code to perform local structural identifiability analysis
Table 3: Local structural identifiability result of the 4-compartmental brain model
Parameter Case1 Case2 Case3 Case4 Case5
Vbb × × ×
Vbm × ×
Vccsf × ×
Vscsf × ×
Qbrain × fixed fixed fixed fixed
Qbulkbc × fixed fixed fixed fixed
Qbulkcb × fixed fixed fixed fixed
QCsink × fixed fixed fixed fixed
QSsink ×
QSin × × fixed fixed fixed
QSout × × fixed fixed fixed
Qglybm × × fixed fixed fixed
Qglyccsf × × fixed fixed fixed
Qglyscsf × × × fixed fixed
PSB × × × fixed fixed
PSC × × × fixed fixed
PSE × × × fixed fixed
CLeffbm × × × × fixed
CLupbm × × × × fixed
CLeffccsf × × × × fixed
CLupccsf × × × × fixed
fubb × × × ×
fubm × × × ×
fuccsf × × × ×
lambb × × × ×
lambm × × × ×
lamccsf × × × ×
Due to the nonlinear nature of the model, parameter identifiability is sometimes challenging. To improve identifiability, we systematically fixed certain parameters, which allowed more parameters to become locally identifiable, as demonstrated in Cases 1–5 of Table (3). In all five cases, data were provided for every state variable.
Practical identifiability analysis
Structural identifiability examines only the mathematical structure of a system of ODEs, the model inputs, and the outputs that are measured. It does not take into account how these outputs are actually measured or the experimental data itself. Practical identifiability analysis, on the other hand, evaluates whether the model parameters can be identified, at least locally, using both the system of ODEs and the available experimental data. It is important to note that a model must be structurally identifiable before it can be practically identifiable. Although several approaches exist for conducting practical identifiability analysis [39, 40], this study focuses specifically on the bootstrapping method. We utilize MONOLIX software to perform bootstrap analysis and present the correlation matrix of the parameters and the confidence intervals for estimated parameters in Example 2.
@@ 4.1 A Step-by-Step Guide to Implementing a Four-Compartment Brain PINN in DeepXDE
    Define exogenous input function: The plasma concentration data are provided as input to the system via the function $g_{1}(t;\theta)$, as defined in Equation (2.6). Since the plasma time - concentration data are available at discrete time points, linear interpolation is applied to construct a continuous approximation. As a result, $g(\theta,t)$ becomes a continuous function of time. To this end we implemented following function (PlasmaInterp) to our program where ”observe t” and ”plasma” are user provided two discrete vectors. It is also worth noting that in practice plasma concentration data are also measured with errors.
@@ 5 Numerical Results
This section presents our main results and their validation. In Example 1, we identify the optimal configuration including the activation function, number of layers and neurons, weight and bias initializers (Glorot uniform and Glorot normal), optimizer, and learning rate that minimizes the total loss (objective function). In Example 2, we solve the inverse problem, estimating the parameters $V_{\mathrm{bb}}$, $V_{\mathrm{bm}}$, $V_{\mathrm{ccsf}}$, $V_{\mathrm{scsf}}$, $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ and presenting the corresponding predicted concentration profiles. We further present the results of the practical identifiability analysis. Finally, in Example 3, we validate these results using two additional numerical and statistical techniques.
@@
As shown in Table (4), the Tanh and Sigmoid activation functions generally perform best for the four-compartment brain model. Figure (8) further confirms their accuracy compared to Sin and ReLU by overlaying observed data with the predicted concentration profiles from each activation function. The training times in Table (4) align with the expected trend that computational cost increases with the number of layers and neurons. However, we observed several occasions where the order of the error (total loss) is of the magnitude of $10^{-2}$ after 10000 iterations. Therefore, we further investigate the best network architecture in Example 2.
@@
Sine 5.69e-02 (571) 3.16e-02 (573) 4.24e-02 (724) 4.57e-02 (1393)
Example 2. In this example, we consider estimating parameters up to six parameters: $V_{\mathrm{bb}}$, $V_{\mathrm{bm}}$, $V_{\mathrm{ccsf}}$, $V_{\mathrm{scsf}}$, $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ though any other parameter set could also be estimated if they are structurally identifiable as shown in the case 5 in Table (3). However, selecting a large number of parameters comes at the cost of increased computational expense. First, we present the results of the practical identification analysis. The correlation matrix derived from the Fisher Information Matrix as shown in Table (5) indicates that most parameter pairs exhibit low to moderate correlations. In particular, the absolute values of the majority of correlation coefficients are below 0.26, suggesting weak linear dependence between parameters and good practical identifiability. The highest correlation is observed between $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ (0.4646), which indicates a moderate positive association but is not sufficiently strong to suggest severe practical non-identifiability. The high correlation between $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ compared to the other parameters is able to explain the high absolute errors of $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ as shown in the Table (7). Overall, the absence of strong correlations (e.g., $|r|>0.8$ ) implies that the selected parameters can be estimated with acceptable reliability from the available data.
Table 5: Correlation matrix of estimated parameters
Vbb Vbm Vccsf Vscsf fubb lamccsf
Vbb 1
Vbm -0.2542 1
Vccsf 0.05475 -0.1292 1
Vscsf 0.05873 -0.02423 -0.151 1
fubb -0.1456 0.1764 -0.009497 -0.1639 1
lamccsf -0.02003 -0.08028 0.1277 -0.01126 0.4646 1
From Table (4) we found that there are several occasions that the error (total loss) is of the magnitude of $10^{-2}$ after 10000 iterations. Among these cases we selected ”tanh” activation function with Glorot normal initializer for initial weight and bias generator. The selected neural network archtechchor contains 6 layers and 50 neurons per layers since this selection shows the closer approximation to the reference parameter values as summarized in the Table (6) after 10000 iterations.
Table 6: Selection of the best neural network architecture after 10,000 iterations
Case 1: (6, 50, tanh) Case 2: (6, 18, sine) Case 3: (1, 27, sigmoid)
Vbb 0.042984428107738494 0.013211678266525268 0.022292330414056777
Vbm 0.704248666763305664 0.617162728309631347 0.730870246887207031
Vccsf 0.093138557821512222 0.057842567563056945 0.060063321143388748
Vscsf 0.015588552691042423 0.010563577413558959 0.011213994637131690
fubb 0.091953142285346984 0.063309334218502044 0.072771822810173034
lamccsf 0.019866755947470664 0.011166308075189590 0.014584582298994064
@@
To further improve the solution we then trained the network for 5 million iterations, which took 54977.47 seconds to complete. After training the model with adam’s optimizer with the learning rate of 0.0001 we further train the model with Limited-memory Broyden-Fletcher-Goldfarb-Shanno optimizer for further smoothness and the best model is found at the 5000018 iteration with total loss of 1.23e051.23e-05. As evidence in Table (7) it can be seen that the parameters $V_{\mathrm{bb}}$, $V_{\mathrm{bm}}$, $V_{\mathrm{ccsf}}$ and $V_{\mathrm{scsf}}$ were trained with very high accuracy. The estimated parameter values of $fu_{\mathrm{bb}}$, and $\lambda_{\mathrm{ccsf}}$ are physiologically acceptable and accurate enough. However, based on the level of accuracy that we expect the model can be further trained.
@@
Table 7: Comparison of reference parameter values, the model estimated parameter values, absolute errors, and 95% confidence intervals (CIs).
Parameter Reference Value PINN Value Absolute Error 95% CI
$V_{bb}$ 0.064952435 0.064952425 $9.34\times 10^{-9}$ $[0.064816,0.065088]$
$V_{bm}$ 1.104115461 1.104115366 $9.41\times 10^{-8}$ $[1.104095,1.104135]$
$V_{ccsf}$ 0.103984624 0.103984609 $1.48\times 10^{-8}$ $[0.103723,0.104246]$
$V_{scsf}$ 0.025996156 0.025996146 $9.28\times 10^{-9}$ $[0.025931,0.026062]$
$fu_{bb}$ 0.125 0.128736447 $3.74\times 10^{-3}$ $[0.111186,0.146286]$
$\lambda_{ccsf}$ 0.026 0.021295592 $4.70\times 10^{-3}$ $[0.012936,0.029656]$
In terms of improving the results through the loss weights, we manually, set the weights (instead of letting network to learn the loss weights) for $\lambda_{IC}$, $\lambda_{ODE}$ and $\lambda_{Data}$ by observing the predictions and comparing the loss curves. In this way, we found that $\lambda_{IC}=(1,1,1,1)$, $\lambda_{ODE}=(2,2,2,2)$, and $\lambda_{Data}=(3,3,3,3)$ yielded the best results in our experiments.
@@
Figure 10: Minimization of weighted log(Loss) components over epoch.
@@
Example 3. In this example, we compare the prediction accuracy of our PBPK-iPINN approach by comparing it with traditional statistical and numerical methods. First, the Stochastic Approximation Expectation-Maximization (SAEM) algorithm [30], implemented in MONOLIX, was used to estimate model parameters. The forward problem was then solved using explicit Runge-Kutta (RK) schemes. Second, the Differential Evolution (DE) algorithm [31], implemented in R, was used for parameter estimation. The forward problem was subsequently solved using the Livermore Solver for Ordinary Differential Equations with Adaptive step-size control, which automatically switches between a non-stiff method (Adams) and a stiff method (Backward Differentiation Formula, BDF) depending on problem stiffness. Physiologically meaningful parameter bounds were applied across all methods to constrain the parameter space and accelerate convergence. For a fair comparison, identical parameter bounds were defined for the PBPK-iPINN, SAEM, and DE algorithms. Absolute errors (for example, $|\text{PINN}_{\text{err}}|=|\text{Reference Value}-\text{PINN}|$), reported in Table (8), were calculated relative to the reference parameter values and rounded to five decimal places.
The results of the SAEM algorithm (implemented in MONOLIX) were obtained after 2000 iterations with 1000 Monte Carlo importance sampling steps. The simulation took 3421 seconds to complete, with the stopping criterion set to a tolerance of $1\times 10^{-8}$. Similarly, the results of the DE algorithm (implemented in R) were obtained after 2000 iterations, with the relative tolerance set to $1\times 10^{-8}$, and required 4083 seconds to complete. All simulations were performed on a MacBook Air (M1, 2020) with 8 GB RAM. The PINN simulation was completed as described in Example 2. Although PINN is computationally expensive, PBPK-iPINN demonstrates promising potential as an alternative to traditional numerical and statistical methods for addressing complex inverse problems, particularly those involving highly stiff systems.